# Varformer: Adapting VAR’s Generative Prior for Image Restoration

### v1

@@ 1 Introduction
Image restoration (IR) aims to reconstruct a high-quality (HQ) image from its degraded low-quality (LQ) counterpart, making it widely applicable in various real-world scenarios, including photo processing, autonomous driving, and surveillance. Recent advances in deep learning have led to powerful IR approaches that excel at addressing specific types of degradation, such as denoising [47, 61, 62], deblurring [37, 38, 40], and low-light enhancement [12, 49, 50], etc. However, these task-specific models struggle with the varied and unpredictable degradations commonly present in real-world scenarios, emphasizing the need for a generalist approach capable of restoring multiple types of degradation.
@@
To alleviate these limitations, some studies has focused on the impact of generative priors within the image restoration (IR) community. These investigations are grounded in the characteristic that generative models inherently learn high-quality image reconstruction, enabling them to maintain robust high-quality priors that facilitate the transition of degraded features to the clean domain. These approaches perform IR tasks using the generative prior [18, 25, 45] from the GAN or diffusion models, achieving superior performance and demonstrating the effectiveness of generative prior for IR tasks. Recently, a novel generative paradigm, VAR, stands out in the image generation community, recognized for its efficiency and generation capabilities that surpass those of Stable Diffusion [39] . VAR generates images with progressively increasing resolutions by utilizing marker mapping for next-scale predictions, resolving the conflict between the inherent bidirectional and two-dimensional structural correlations in image patch labeling and the unidirectional nature of the autoregressive model.
Motivated by the powerful generative capabilities of VAR and the alignment between its endogenous scale representation and the widely recognized multi-scale restoration principle, we aim to explore the rich prior knowledge embedded in the pre-trained VAR across the scale dimension for restoration tasks. Our investigation reveals the adaptability of the autoregressive scale representation, as the VAR feature representation shifts from capturing wide-ranging global patterns at lower scales to highlighting intricate details at higher scales. As described in Fig. 2 a and b, we apply pre-trained VAR to reconstruct the paired clean and degraded images, respectively. When we replace the partially scale features of the degraded images obtained from the VQVAE encoder with the autoregressive generated scale features, the degradation of the images gradually diminishes. Specifically, Fig. 2 c shows that substituting lower scales can alleviate global degradations (e.g., low light and haze), while higher scales address local degradations (e.g., noise and rain). For a boarder investigation, we further perform a statistical analysis of the prior knowledge related to high-quality images, utilizing 100 paired images for each type of degradation. As illustrated in Fig. 3 , VAR’s next-scale prediction projects the multi-scale latent representations of both degraded and clean images into a common space, thereby reducing the distribution gap between various degradations and clean images, demonstrating the model’s adaptive distribution alignment capability.
@@ 2.1 Image Restoration
The purpose of image restoration is to reconstruct high quality natural images from the degraded images (e.g. noise, blur, rain drops). Early methods typically focuses on incorporating various natural image priors along with hand-crafted features for specific degradation removal tasks [4, 13]. Recently, deep learning based methods have made compelling progress on various image restoration tasks. For instance, DGUNet is proposed based on Proximal Gradient Descent (PGD) algorithm for a gradient estimation strategy without loss of interpretability [34]. IDR [60] employs an ingredients-oriented paradigm to investigate the correlation among various restoration tasks. TransWeather [42] designs a transformer-based network with learnable weather type queries to tackle various weather degradation.
@@ 3.1 Generative Priors in VAR for IR
Different from traditional autoregressive methods, VAR introduces a novel visual autoregressive modeling paradigm, shifting from ”next-token prediction” to ”next-scale prediction”. This paradigm can solve mathematical inconsistencies and structural degradation, which is more optimal for generating highly-structured images. In VAR, each unit predicts an entire token map at a different scale. Starting with a 1 × 1 token map ${r}_{1}$, VAR predicts a sequence of multi-scale token maps $({r}_{1},{r}_{2},\dots,{r}_{K})$, increasing in resolution. The generation process is expressed as:
$p({r}_{1},{r}_{2},\dots,{r}_{K})=\displaystyle\prod_{k=1}^{K}p({r}_{k}|{r}_{1},{r}_{2},\dots,{r}_{k-1})$ (1)
@@
${I}_{rec}=D(\displaystyle\sum_{i\notin C}^{K}{F}_{{e}_{i}}^{deg}+\displaystyle\sum_{i\in C}^{K}{F}_{{v}_{i}}^{deg})$ (2)
@@ 3.2 Architecture of the VarFormer
Adapter for Domain Shift (Train in Stage 1). Due to the existence of distributional shift between high quality sources for VAR pre-training and degraded datasets for image restoration, direct reuse of the features from the Encoder of VAR may be suboptimal. However, fine-tuning disrupts the the knowledge in VAR. Therefore, inspired by AWRCP [53], we chose to freeze the encoder and insert an Adapter containing self-attention blocks after it to retain pre-trained knowledge while narrowing the domain gap.
@@
${w}_{1}^{g},{w}_{2}^{g}=Softmax(Conv(RSTBs(Concat({F}_{e/d}^{i},\widehat{S}_{w}^{i}))))$
${F}_{{g}_{e/d}}^{i}={F}_{e/d}^{i}\times{w}_{1}^{g}+\widehat{S}_{w}^{i}\times{w}_{2}^{g}$ (4)
@@
${F}_{in}^{i+1}={A}_{q,m}\cdot({A}_{m,k}\cdot V)+{F}_{e/d}^{i}$
${A}_{q,m}=Softmax(Q\cdot{M}^{T}/\sqrt{d})$ (5)
${A}_{m,k}=Softmax(M\cdot{K}^{T}/\sqrt{d})$
@@
${F}_{d}^{i}=\sigma({\theta}_{i})\times\widehat{F}_{d}^{i}+(1-\sigma({\theta}_{i}))\times{F}_{e}^{i}$ (6)
@@
Figure 6: The visualization results of zero-shot generalization in real-world TOLED dataset. Please zoom in for details.
@@ 3.3 Training Objectives
${\mathcal{L}}_{fema}=\displaystyle\sum_{i=1}^{K}-{s}_{i}log(\widehat{s}_{i})+{\left\|{F}_{a}-sg({F}_{{e}_{gt}}^{q})\right\|}_{2}^{2}$ (7)
@@
${\mathcal{L}}_{rec}={-PSNR({I}_{gt},{I}_{rec})+{\left\|\psi({I}_{gt})-\psi({I}_{rec})\right\|}_{2}^{2}}$ (8)
@@ 4 Experiments
Table 1: Quantitative comparison with state-of-the-art task-specific methods and universal methods on image deraining, low-light image enhancement, motion deblurring and image dehazing tasks.
@@
IDR [60] 2023 30.87 0.906 27.94 0.848 22.32 0.836 25.33 0.945
DA-CLIP [30] 2023 28.75 0.844 26.24 0.801 24.27 0.885 31.42 0.941
Prompt-IR [36] 2023 29.44 0.848 27.31 0.834 22.95 0.844 32.17 0.953
DiffUIR [64] 2024 31.14 0.907 29.88 0.874 25.02 0.901 32.74 0.944
@@
AirNet [21] 33.49 30.91 27.66 33.16 30.83 27.45 34.14 31.74 28.59
IDR [60] 34.05 31.67 28.05 32.92 31.29 28.45 34.53 32.22 28.93
DA-CLIP [30] 30.16 28.89 27.04 33.14 30.99 27.61 33.86 32.30 28.84
Prompt-IR [36] 32.17 29.89 28.03 33.04 31.89 27.72 33.78 32.21 28.64
DiffUIR [64] 33.86 30.88 26.63 32.19 29.65 25.87 33.24 30.70 27.19
@@
AirNet [21] 38.32 0.945
Prompt-IR [36] 39.52 0.954
DA-CLIP [30] 34.04 0.824
IDR [60] 39.74 0957
MPerceiver [3] 39.96 0.959
@@ 4.1 Experimental Settings
Implementation Details. We adopt the Adam optimizer [19](${\beta}_{1}=0.9$, ${\beta}_{2}=0.999$) with the initial learning rate 1e-4 gradually reduced to 1e-6 with cosine annealing to train our model.We random crop $256\times 256$ patch from original image as network input after data augumentation. As the data size varies greatly from task to task, we set the weight of each task in one batch as 0.3 for dehazing, 0.1 for low-light, 0.2 for deraining, 0.2 for Gaussian denoising,0.1 for real image denoising, and 0.1 for motion deblurring.
@@ 4.2 All-in-One Restoration
We compare VarFormer with both task-specific and universal methods. To ensure a fair evaluation, we train all-in-one models from scratch using our training strategy. The quantitative results are presented in  Tabs. 1, 2 and 3, where it can be observed that VarFormer demonstrates superior performance across all 6 tasks. Notably, in low-light enhancement and dehazing tasks, VarFormer’s performance with the all-in-one training strategy even surpasses that of task-specific methods, highlighting its exceptional capabilities. We also provide visual comparisons with other state-of-the-art universal methods, as shown in Fig. 5, with more results available in the supplementary material. It is evident that, compared to other universal methods, VarFormer yields more steady results in all image restoration tasks.
@@ 4.3 Zero-shot Results
In Tab. 4, we present the generalization performance of various methods on unseen tasks without fine-tuning. Specifically, high-resolution images captured under under-display camera (UDC) systems often suffer from various types of degradations due to the point spread function and lower light transmission rate, posing challenges for the prediction of restoration models. Encouragingly, our VarFormer achieves state-of-the-art performance. For instance, as shown in Fig. 6, compared to other universal methods, our approach effectively addresses blurring.
@@
4.4 Abalation Studies
In this section, we perform a series of ablation studies to better the effectiveness of our designs. In Tab. 5, we conducte ablation studies on the design of key components, including the adaptive mix-up skip that integrates encoder features into the decoder, the adapter that addresses the domain shift issue when using pre-trained VAR on degraded datasets, and the DAE and AFT modules responsible for generating and fusing multi-scale distribution alignment priors. We evaluate the performance based on the average performance across six tasks. The results indicate a gradual improvement with the addition of each component and a corresponding decline when each component is removed, underscoring the effectiveness of each module.
Figure 7: The training trajectory using different number of DAE.
@@ 4.4 Abalation Studies
Table 6: Ablation experiments on the DAE number.
Exp. DAE PSNR↑ SSIM↑
all skip1 skip2 w/o
a 28.98 0.867
b 29.29 0.883
c 29.49 0.902
d 29.66 0.926
Additionally, we explored the influence of the guidance strength of VAR multi-scale priors on restoration tasks by altering the quantity of DAE modules. The results from Tab. 6 indicate a positive correlation between model performance and the guidance strength provided by VAR priors. This discovery indicates that the multi-scale priors within VAR exhibit notable effectiveness in directing the optimization trajectory of intricate multi-task restoration endeavors, effectively steering the model towards enhancing performance. The training trajectories illustrated in Fig. 7 also support this conclusion, indicating that the integration of DAE modules can accelerate the model’s convergence.

### v2

# Navigating Image Restoration with VAR’s Distribution Alignment Prior

@@ 1 Introduction
Image restoration (IR) aims to reconstruct a high-quality (HQ) image from its degraded low-quality (LQ) counterpart, making it widely applicable in various real-world scenarios, including photo processing, autonomous driving, surveillance, segmentation and detection  [8, 16, 7]. Recent advances in deep learning have led to powerful IR approaches that excel at addressing specific types of degradation, such as denoising [48, 62, 63], deblurring [38, 39, 41], and low-light enhancement [13, 50, 51], etc. However, these task-specific models struggle with varied and unpredictable degradations in real-world scenarios. This highlights the need for a generalist approach capable of addressing multiple degradation types.
@@
To alleviate these limitations, some studies have focused on the impact of generative priors within the image restoration (IR) community. These investigations are grounded in the characteristic that generative models inherently learn high-quality image reconstruction, enabling them to maintain robust high-quality priors that facilitate the transition of degraded features to the clean domain. These approaches perform IR tasks using the generative prior [20, 27, 46] from the GAN or diffusion models, achieving superior performance and demonstrating the effectiveness of generative prior for IR tasks. Recently, a novel generative paradigm, VAR, stands out in the image generation community, recognized for its efficiency and generation capabilities that surpass those of Stable Diffusion [40] . VAR generates images with progressively increasing resolutions by utilizing marker mapping for next-scale predictions, resolving the conflict between the inherent bidirectional and two-dimensional structural correlations in image patch labeling and the unidirectional nature of the autoregressive model.
Motivated by the powerful generative capabilities of VAR and the alignment between its endogenous scale representation and the widely recognized multi-scale restoration principle, we aim to explore the rich prior knowledge embedded in the pre-trained VAR across the scale dimension for restoration tasks. Our investigation reveals the adaptability of the autoregressive scale representation, as the VAR feature representation shifts from capturing wide-ranging global patterns at lower scales to highlighting intricate details at higher scales. As described in Fig. 2 (a)-(c), we apply pre-trained VAR to reconstruct the paired clean and degraded images, respectively. When we replace the partially scale features of the degraded images obtained from the VQVAE encoder with the autoregressive generated scale features, the degradation of the images gradually diminishes.
For a boarder investigation, we further perform a statistical analysis of the prior knowledge related to high-quality images, utilizing 100 paired images for each type of degradation. As illustrated in Fig. 3, VAR’s next-scale prediction projects the multi-scale latent representations of both degraded and clean images into a common space, thereby reducing the distribution gap between various degradations and clean images, demonstrating the model’s adaptive distribution alignment capability.
@@ 2.1 Image Restoration
The purpose of image restoration is to reconstruct high quality natural images from the degraded images (e.g. noise, blur, rain drops) . Early methods typically focuses on incorporating various natural image priors along with hand-crafted features for specific degradation removal tasks [3, 14]. Recently, deep learning based methods have made compelling progress on various image restoration tasks. For instance, DGUNet [35] is proposed based on Proximal Gradient Descent (PGD) algorithm for a gradient estimation strategy without loss of interpretability [35]. IDR [61] employs an ingredients-oriented paradigm to investigate the correlation among various restoration tasks. TransWeather [43] designs a transformer-based network with learnable weather type queries to tackle various weather degradation.
@@ 3.1 Generative Priors in VAR for IR
Different from traditional autoregressive methods, VAR introduces a novel visual autoregressive modeling paradigm, shifting from “next-token prediction” to “next-scale prediction”. This paradigm can solve mathematical inconsistencies and structural degradation, which is more optimal for generating highly-structured images. In VAR, each unit predicts an entire token map at a different scale. Starting with a 1 × 1 token map ${r}_{1}$, VAR predicts a sequence of multi-scale token maps $({r}_{1},{r}_{2},\dots,{r}_{K})$, increasing in resolution. The generation process is expressed as:
$p({r}_{1},{r}_{2},\dots,{r}_{K})=\displaystyle\prod_{k=1}^{K}p({r}_{k}|{r}_{1},{r}_{2},\dots,{r}_{k-1}),$ (1)
@@
${I}_{rec}=D(\displaystyle\sum_{i\notin C}^{K}{F}_{{e}_{i}}^{deg}+\displaystyle\sum_{i\in C}^{K}{F}_{{v}_{i}}^{deg}),$ (2)
@@ 3.2 Architecture of the VarFormer
Adapter for Domain Shift (Train in Stage 1). Due to the existence of distributional shift between high quality sources for VAR pre-training and degraded datasets for image restoration, direct reuse of the features from the Encoder of VAR may be suboptimal. However, fine-tuning disrupts the knowledge in VAR. Therefore, inspired by AWRCP [54], we chose to freeze the encoder and insert an Adapter containing self-attention blocks after it to retain pre-trained knowledge while narrowing the domain gap.
@@
${w}_{1}^{g},{w}_{2}^{g}=Softmax(Conv(RSTBs(Concat({F}_{e/d}^{i},\widehat{S}_{w}^{i})))),$
${F}_{{g}_{e/d}}^{i}={F}_{e/d}^{i}\times{w}_{1}^{g}+\widehat{S}_{w}^{i}\times{w}_{2}^{g},$ (4)
@@
${F}_{in}^{i+1}={A}_{q,m}\cdot({A}_{m,k}\cdot V)+{F}_{e/d}^{i},$
${A}_{q,m}=Softmax(Q\cdot{M}^{T}/\sqrt{d}),$ (5)
${A}_{m,k}=Softmax(M\cdot{K}^{T}/\sqrt{d}),$
@@
${F}_{d}^{i}=\sigma({\theta}_{i})\times\widehat{F}_{d}^{i}+(1-\sigma({\theta}_{i}))\times{F}_{e}^{i},$ (6)
@@
Figure 6: Visual comparison with state-of-the-art methods on low-light image enhancement task. Please zoom in for details.
@@ 3.3 Training Objectives
${\mathcal{L}}_{fema}=\displaystyle\sum_{i=1}^{K}-{s}_{i}log(\widehat{s}_{i})+{\left\|{F}_{a}-sg({F}_{{e}_{gt}}^{q})\right\|}_{2}^{2},$ (7)
@@
${\mathcal{L}}_{rec}={-PSNR({I}_{gt},{I}_{rec})+{\left\|\psi({I}_{gt})-\psi({I}_{rec})\right\|}_{2}^{2}},$ (8)
@@ 4 Experiments
Table 1: Quantitative comparison with task-specific methods and universal methods on image deraining, low-light image enhancement, motion deblurring and image dehazing tasks. $\dagger$ means reimplementing in our datasets with the same settings for fair comparison.
@@
IDR [61$\dagger$ 2023 30.87 0.906 27.94 0.848 22.32 0.836 25.33 0.945
DA-CLIP [32$\dagger$ 2023 28.75 0.844 26.24 0.801 24.27 0.885 31.42 0.941
Prompt-IR [37$\dagger$ 2023 29.44 0.848 27.31 0.834 22.95 0.844 32.17 0.953
DiffUIR [65$\dagger$ 2024 31.14 0.907 29.88 0.874 25.02 0.901 32.74 0.944
@@
AirNet [23$\dagger$ 33.49 30.91 27.66 33.16 30.83 27.45 34.14 31.74 28.59
IDR [61$\dagger$ 34.05 31.67 28.05 32.92 31.29 28.45 34.53 32.22 28.93
DA-CLIP [32$\dagger$ 30.16 28.89 27.04 33.14 30.99 27.61 33.86 32.30 28.84
Prompt-IR [37$\dagger$ 32.17 29.89 28.03 33.04 31.89 27.72 33.78 32.21 28.64
DiffUIR [65$\dagger$ 33.86 30.88 26.63 32.19 29.65 25.87 33.24 30.70 27.19
@@
AirNet [23$\dagger$ 38.32 0.945
@@
IDR [61$\dagger$ 39.74 0.957
DA-CLIP [32$\dagger$ 34.04 0.824
Prompt-IR [37$\dagger$ 39.52 0.954
DiffUIR [65$\dagger$ 40.11 0.976
@@ 4.1 Experimental Settings
Implementation Details. We adopt the Adam optimizer [21](${\beta}_{1}=0.9$, ${\beta}_{2}=0.999$) with the initial learning rate 1e-4 gradually reduced to 1e-6 with cosine annealing to train our model. We random crop $256\times 256$ patch from original image as network input after data augmentation. As the data size varies greatly from task to task, we set the weight of each task in one batch as 0.3 for dehazing, 0.1 for low-light, 0.2 for deraining, 0.2 for Gaussian denoising,0.1 for real image denoising, and 0.1 for motion deblurring.
@@ 4.2 All-in-One Restoration
We compare VarFormer with both task-specific and universal methods. To ensure a fair evaluation, we train all-in-one models from scratch using our training strategy. The quantitative results are presented in  Tabs. 1, 2 and 3, where it can be observed that VarFormer demonstrates superior performance across all 6 tasks. Notably, in low-light enhancement and dehazing tasks, VarFormer’s performance with the all-in-one training strategy even surpasses that of task-specific methods, highlighting its exceptional capabilities. We also provide visual comparisons with other state-of-the-art universal methods, as shown in Fig. 5 and Fig. 6, with more results available in the supplementary material. It is evident that, compared to other universal methods, VarFormer yields more steady results in all image restoration tasks.
@@ 4.3 Zero-shot Results
In Tab. 4, we present the generalization performance of various methods on unseen tasks without fine-tuning. Specifically, high-resolution images captured under under-display camera (UDC) systems often suffer from various types of degradations due to the point spread function and lower light transmission rate, posing challenges for the prediction of restoration models. Encouragingly, our VarFormer achieves state-of-the-art performance.
@@
4.4 Ablation Studies
In this section, we perform a series of ablation studies to better the effectiveness of our designs. In Tab. 5, we conduct ablation studies on the design of key components, including the adaptive mix-up skip that integrates encoder features into the decoder, the adapter that addresses the domain shift issue when using pre-trained VAR on degraded datasets, and the DAE and AFT modules responsible for generating and fusing multi-scale distribution alignment priors. We evaluate the performance based on the average performance across six tasks. The results indicate a gradual improvement with the addition of each component and a corresponding decline when each component is removed, underscoring the effectiveness of each module.
To verify the sensitivity of DAE module to various types of degradation, We visualize the region-specific fusion weights ${w}_{1}^{g}$ and ${w}_{2}^{g}$ within the DAE module in Fig. 7. Here, weight ${w}_{1}^{g}$ is for the modulated feature map, and weight ${w}_{2}^{g}$ is for the VAR distribution alignment prior. It is observed that the areas highlighted by the VAR distribution alignment prior closely match the regions that have been adversely affected by degradation. This indicates that DAE module can accurately identify the degraded areas and effectively guide the introduction of high-quality distribution alignment priors for the restoration process.
Figure 7: Visualization of weight maps from DAE.
@@ 5 Conclusion
Acknowledgement. This work was supported by the Anhui Provincial Natural Science Foundation under Grant 2108085UD12. We acknowledge the support of GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC.